OpenStudy (anonymous):

Solve each equation. Give answers involving radicals in simplest radical form.

OpenStudy (anonymous):

you getting the questions? :P

OpenStudy (anonymous):

\[z^2 + iz + 2=0\]

OpenStudy (anonymous):

yes ^^ that is the equation...

OpenStudy (anonymous):

z2+iz+2=0 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. z2+iz=−2 Factor the left-hand side of the equation. z(z+i)=−2 Set the single term factor on the left-hand side of the equation equal to 0. z=0 Set each of the factors of the left-hand side of the equation equal to 0. z+i=0 Solve the equations for z. z=0,−i

OpenStudy (anonymous):

ez XD

OpenStudy (anonymous):

but the answer is {-2i,i} in the answer key...

OpenStudy (anonymous):

I don't know how they got that answer.

OpenStudy (anonymous):

yea... i don't either

OpenStudy (anonymous):

I figured it out on my own! yaeah me! I deserve a medal!